# Bounding the SNPR distance between two tree-child networks using generalised agreement forests

### v1

@@ 6. Bounding the tree-child SNPR distance
    Let $(u,v)$ be a reticulation edge of ${\mathcal{N}}$ such that $(u,v)\notin{\mathcal{R}}$. If ${\mathcal{N}}^{\prime}$ can be obtained from ${\mathcal{N}}$ by an SNPR- that deletes $(u,v)$ and suppresses $u$ and $v$, then ${\mathcal{D}}$ is a tree-child digraph of ${\mathcal{N}}^{\prime}$ and there exists an extension ${\mathcal{R}}^{\prime}$ of ${\mathcal{D}}$ in ${\mathcal{N}}^{\prime}$ such that
    $|E_{{\mathcal{N}}}-E_{{\mathcal{R}}}|-1=|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|.$
@@
    Let $e=(u,w)$ and $e^{\prime}=(u^{\prime},w^{\prime})$ be two distinct tree edges of ${\mathcal{N}}$. If ${\mathcal{N}}^{\prime}$ can be obtained from ${\mathcal{N}}$ by an SNPR+ that subdivides $e$ and $e^{\prime}$ with a new vertex $v$ and $v^{\prime}$, respectively and adds the new reticulation edge $(v^{\prime},v)$, then ${\mathcal{D}}$ is a tree-child digraph of ${\mathcal{N}}^{\prime}$ and there exists an extension ${\mathcal{R}}^{\prime}$ of ${\mathcal{D}}$ in ${\mathcal{N}}^{\prime}$ such that
    $|E_{{\mathcal{N}}}-E_{{\mathcal{R}}}|+1=|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|.$
@@
    Let $e=(u,v)$ be a tree edge of ${\mathcal{N}}$ such that $e\notin{\mathcal{R}}$, and let $f=(p_{u^{\prime}},c_{u^{\prime}})$ be an edge in ${\mathcal{N}}$ that is distinct from $e$. If ${\mathcal{N}}^{\prime}$ can be obtained from ${\mathcal{N}}$ by an SNPR± that deletes $e$, suppresses $u$, subdivides $f$ with a new vertex $u^{\prime}$, and adds the new edge $(u^{\prime},v)$, then ${\mathcal{D}}$ is a tree-child digraph of ${\mathcal{N}}^{\prime}$ and there exists an extension ${\mathcal{R}}^{\prime}$ of ${\mathcal{D}}$ in ${\mathcal{N}}^{\prime}$ such that
    $|E_{{\mathcal{N}}}-E_{{\mathcal{R}}}|=|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|.$
@@
We first establish (i). Since $(u,v)\notin{\mathcal{R}}$ and ${\mathcal{D}}$ is a tree-child digraph of ${\mathcal{N}}$, it follows that ${\mathcal{D}}$ is also such a digraph of ${\mathcal{N}}^{\prime}$. Now, let ${\mathcal{R}}^{\prime}$ be the digraph obtained from ${\mathcal{R}}$ by applying the following operation to each vertex $w\in\{u,v\}$. If $w$ has in-degree zero and out-degree one in ${\mathcal{R}}$, then delete $w$ and, if $w$ has in-degree one and out-degree one in ${\mathcal{R}}$, then suppress $w$. As $(u,v)\not\in{\mathcal{R}}$, it follows by Lemma 4.1 that each of $u$ and $v$ has either in-degree zero and out-degree one, or in-degree one and out-degree one in ${\mathcal{R}}$. Thus, ${\mathcal{R}}^{\prime}$ is well defined. It now follows from the construction that ${\mathcal{R}}^{\prime}$ is an extension of ${\mathcal{D}}$ in ${\mathcal{N}}^{\prime}$ with $|E_{{\mathcal{N}}}-E_{{\mathcal{R}}}|-1=|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|$.
@@
We now turn to (iii). Let $p_{u}$ be the parent of $u$, and let $c_{u}$ be the child of $u$ that is not $v$. Since $u$ is a tree vertex, $p_{u}$ and $c_{u}$ are well defined. Furthermore, let ${\mathcal{M}}$ be the embedding of ${\mathcal{D}}$ in ${\mathcal{N}}$ that underlies ${\mathcal{R}}$. By Lemma 4.1(i), $(u,c_{u})\in{\mathcal{R}}$. If $(p_{u},u)\notin{\mathcal{R}}$, then the degree constraints of the vertices in ${\mathcal{D}}$ imply that $(u,c_{u})\notin{\mathcal{M}}$. On the other hand, if $(p_{u},u)\in{\mathcal{R}}$, then $(u,c_{u})\in{\mathcal{R}}$ and, in turn, either both of $(p_{u},u)$ and $(u,c_{u})$ are in ${\mathcal{M}}$ or neither. Now obtain ${\mathcal{M}}^{\prime}$ from ${\mathcal{M}}$ by replacing $(p_{u},u)$ and $(u,c_{u})$ with $(p_{u},c_{u})$ if $(p_{u},u)$ and $(u,c_{u})$ are in ${\mathcal{M}}$, and replacing $f$ with $(p_{u^{\prime}},u^{\prime})$ and $(u^{\prime},c_{u^{\prime}})$ if $f\in{\mathcal{M}}$. Thus, as ${\mathcal{M}}$ is the embedding of ${\mathcal{D}}$ in ${\mathcal{N}}$ that underlies ${\mathcal{R}}$, it follows from the construction of ${\mathcal{N}}^{\prime}$ from ${\mathcal{N}}$ that ${\mathcal{M}}^{\prime}$ is also an embedding of ${\mathcal{D}}$ in ${\mathcal{N}}^{\prime}$. Hence, ${\mathcal{D}}$ is a tree-child digraph of ${\mathcal{N}}^{\prime}$. To complete the proof, let ${\mathcal{R}}^{\prime}$ be the digraph obtained from ${\mathcal{R}}$ by applying the following operations. If $u$ has in-degree zero in ${\mathcal{R}}$, then delete $u$ and, if $u$ has in-degree one in ${\mathcal{R}}$, then suppress $u$. Moreover, if $f\notin{\mathcal{R}}$, then add $(u^{\prime},c_{u^{\prime}})$ and, if $f\in{\mathcal{R}}$, then subdivide $(p_{u^{\prime}},c_{u^{\prime}})$ with $u^{\prime}$. Again, by Lemma 4.1, $u$ has either in-degree zero and out-degree one, or in-degree one and out-degree one in ${\mathcal{R}}$ and, so, ${\mathcal{R}}^{\prime}$ is well defined. As ${\mathcal{R}}$ is an extension of ${\mathcal{D}}$ in ${\mathcal{N}}$ with $e\notin{\mathcal{R}}$ and $(u^{\prime},v)\notin{\mathcal{R}}^{\prime}$, it now follows that ${\mathcal{R}}^{\prime}$ is an extension of ${\mathcal{D}}$ in ${\mathcal{N}}^{\prime}$ with $|E_{{\mathcal{N}}}-E_{{\mathcal{R}}}|=|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|$. ∎
The next lemma and corollary show that there always exists a tree-child $\mathrm{SNPR}$ sequence connecting two tree-child networks with certain desirable properties. These properties will be leveraged later to establish one of the two inequalities of Theorem 1.1.
Lemma 6.2.
@@
    $({\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{i},{\mathcal{N}}_{i+3},{\mathcal{N}}_{i+4},\ldots,{\mathcal{N}}_{t-1},{\mathcal{N}}_{t})$ is a tree-child $\mathrm{SNPR}$ sequence connecting ${\mathcal{N}}_{0}$ and ${\mathcal{N}}_{t}$ of length $t-2$,
@@
    $({\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{i},{\mathcal{N}}_{i+2},{\mathcal{N}}_{i+3},\ldots,{\mathcal{N}}_{t-1},{\mathcal{N}}_{t})$ is a tree-child $\mathrm{SNPR}$ sequence connecting ${\mathcal{N}}_{0}$ and ${\mathcal{N}}_{t}$ of length $t-1$ such that ${\mathcal{N}}_{i+2}$ is obtained from ${\mathcal{N}}_{i}$ by an $\mathrm{SNPR}^{\pm}$,
@@
    $({\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{i},{\mathcal{N}}_{i+2},{\mathcal{N}}_{i+3},\ldots,{\mathcal{N}}_{t-1},{\mathcal{N}}_{t})$ is a tree-child $\mathrm{SNPR}$ sequence connecting ${\mathcal{N}}_{0}$ and ${\mathcal{N}}_{t}$ of length $t-1$ such that ${\mathcal{N}}_{i+2}$ is obtained from ${\mathcal{N}}_{i}$ by an $\mathrm{SNPR}^{-}$, or
@@
    there exists a tree-child $\mathrm{SNPR}$ sequence
    $({\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{i},{\mathcal{N}}_{i+1}^{\prime},{\mathcal{N}}_{i+2},{\mathcal{N}}_{i+3},\ldots,{\mathcal{N}}_{t-1},{\mathcal{N}}_{t})$
    connecting ${\mathcal{N}}_{0}$ and ${\mathcal{N}}_{t}$ of length $t$ such that ${\mathcal{N}}_{i+1}^{\prime}$ is obtained from ${\mathcal{N}}_{i}$ by an $\mathrm{SNPR}^{-}$ and ${\mathcal{N}}_{i+2}$ is obtained from ${\mathcal{N}}_{i+1}^{\prime}$ by an $\mathrm{SNPR}^{+}$ (resp. $\mathrm{SNPR}^{\pm}$).
@@
Suppose that there exists an element $i\in\{0,1,2,\ldots,t-2\}$ such that ${\mathcal{N}}_{i+1}$ is obtained from ${\mathcal{N}}_{i}$ by an SNPR+ and ${\mathcal{N}}_{i+2}$ is obtained from ${\mathcal{N}}_{i+1}$ by an SNPR-. Let $e=(u,v)$ be the reticulation edge of ${\mathcal{N}}_{i+1}$ that is added in obtaining ${\mathcal{N}}_{i+1}$ from ${\mathcal{N}}_{i}$. Furthermore, let $e^{\prime}=(u^{\prime},v^{\prime})$ be the reticulation edge in ${\mathcal{N}}_{i+1}$ that is deleted in obtaining ${\mathcal{N}}_{i+2}$ from ${\mathcal{N}}_{i+1}$. If $e=e^{\prime}$, then ${\mathcal{N}}_{i}\cong{\mathcal{N}}_{i+2}$, and so (i) holds. We may therefore assume that $e\neq e^{\prime}$. If $v=v^{\prime}$, let $w$ be the child of $v$ in ${\mathcal{N}}_{i+1}$, and let $c_{u}$ and $p_{u}$ be the child and parent, respectively, of $u$ in ${\mathcal{N}}_{i+1}$ such that $c_{u}\neq v$. Since $u$ is a tree vertex, $c_{u}$ and $p_{u}$ are well defined. Observe that $(u^{\prime},w)$ and $(p_{u},c_{u})$ are tree edges in ${\mathcal{N}}_{i}$. It now follows that ${\mathcal{N}}_{i+2}$ can be obtained from ${\mathcal{N}}_{i}$ by the SNPR± that deletes $(u^{\prime},w)$, suppresses $u^{\prime}$, subdivides $(p_{u},c_{u})$ with a new vertex $u$, and adds the edge $(u,w)$, and so (ii) holds. So assume that $e\neq e^{\prime}$ and $v\neq v^{\prime}$. As ${\mathcal{N}}_{i+1}$ is tree-child, it follows that $e^{\prime}$ is not incident with $u$ or $v$. Thus, $e^{\prime}$ is a reticulation edge of ${\mathcal{N}}_{i}$. Let ${\mathcal{N}}_{i+1}^{\prime}$ be the phylogenetic network obtained from ${\mathcal{N}}_{i}$ by deleting $e^{\prime}$ and suppressing $u^{\prime}$ and $v^{\prime}$. By Lemma 2.2, ${\mathcal{N}}_{i+1}^{\prime}$ is tree-child. It is now straightforward to check that ${\mathcal{N}}_{i+2}$ can be obtained from ${\mathcal{N}}_{i+1}^{\prime}$ by an SNPR+, and so (iv) holds.

### v2

@@ 6 Bounding the tree-child SNPR distance
    Let $(u,v)$ be a reticulation edge of $\mathcal{N}$} such that $(u,v)\notin{\mathcal{R}}$. If $\mathcal{N}$$^{\prime}$ can be obtained from $\mathcal{N}$} by an SNPR- that deletes $(u,v)$ and suppresses $u$ and $v$, then $\mathcal{D}$} is a tree-child digraph of $\mathcal{N}$$^{\prime}$ and there exists an extension $\mathcal{R}$$^{\prime}$ of $\mathcal{D}$} in $\mathcal{N}$$^{\prime}$ such that —E_N-E_R—-1=—E_N’-E_R’—.
@@
    Let $e=(u,w)$ and $e^{\prime}=(u^{\prime},w^{\prime})$ be two distinct tree edges of $\mathcal{N}$}. If $\mathcal{N}$$^{\prime}$ can be obtained from $\mathcal{N}$} by an SNPR+ that subdivides $e$ and $e^{\prime}$ with a new vertex $v$ and $v^{\prime}$, respectively, and adds the new reticulation edge $(v^{\prime},v)$, then $\mathcal{D}$} is a tree-child digraph of $\mathcal{N}$$^{\prime}$ and there exists an extension $\mathcal{R}$$^{\prime}$ of $\mathcal{D}$} in $\mathcal{N}$$^{\prime}$ such that —E_N-E_R—+1=—E_N’-E_R’—.
@@
    Let $e=(u,v)$ be a tree edge of $\mathcal{N}$} such that $e\notin{\mathcal{R}}$, and let $f=(p_{u^{\prime}},c_{u^{\prime}})$ be an edge in $\mathcal{N}$} that is distinct from $e$. If $\mathcal{N}$$^{\prime}$ can be obtained from $\mathcal{N}$} by an SNPR± that deletes $e$, suppresses $u$, subdivides $f$ with a new vertex $u^{\prime}$, and adds the new edge $(u^{\prime},v)$, then $\mathcal{D}$} is a tree-child digraph of $\mathcal{N}$$^{\prime}$ and there exists an extension $\mathcal{R}$$^{\prime}$ of $\mathcal{D}$} in $\mathcal{N}$$^{\prime}$ such that —E_N-E_R—=—E_N’-E_R’—.
[…]
Figure 6: Setup of parts (i) and (iii) in the proof of Lemma 10. Red lines indicate vertices and edges in $\mathcal{R}$} and $\mathcal{R}$$^{\prime}$. In (i), observe that $u$ has in-degree zero and out-degree one, and $v$ has in-degree one and out-degree one in $\mathcal{R}$} whereas, in (iii), observe that $u$ has in-degree one and out-degree one, $(p_{u^{\prime}},c_{u^{\prime}})\notin{\mathcal{R}}$, and $v$ may or may not be a leaf in $\mathcal{N}$}. Vertex labels $y_{1}$, $y_{2}$, $y_{3}$, and $y_{4}$ are only used to clarify the figure.
@@
The setup used to establish parts (i) and (iii) is illustrated in Figure 6. We first establish (i). Since $(u,v)\notin{\mathcal{R}}$ and $\mathcal{D}$} is a tree-child digraph of $\mathcal{N}$}, it follows that $\mathcal{D}$} is also such a digraph of $\mathcal{N}$$^{\prime}$. Now, let $\mathcal{R}$$^{\prime}$ be the digraph obtained from $\mathcal{R}$} by applying the following operation to each vertex $w\in\{u,v\}$. If $w$ has in-degree zero and out-degree one in $\mathcal{R}$}, then delete $w$ and, if $w$ has in-degree one and out-degree one in $\mathcal{R}$}, then suppress $w$. As $(u,v)\not\in{\mathcal{R}}$, it follows by Lemma 5 that each of $u$ and $v$ has either in-degree zero and out-degree one, or in-degree one and out-degree one in $\mathcal{R}$}. Thus, $\mathcal{R}$$^{\prime}$ is well defined. It now follows from the construction that $\mathcal{R}$$^{\prime}$ is an extension of $\mathcal{D}$} in $\mathcal{N}$$^{\prime}$ with $|E_{{\mathcal{N}}}-E_{{\mathcal{R}}}|-1=|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|$.
@@
We now turn to (iii). Let $p_{u}$ be the parent of $u$, and let $c_{u}$ be the child of $u$ that is not $v$. Since $u$ is a tree vertex, $p_{u}$ and $c_{u}$ are well defined. Furthermore, let $\mathcal{M}$} be the embedding of $\mathcal{D}$} in $\mathcal{N}$} that underlies $\mathcal{R}$}. By Lemma 5(i), $(u,c_{u})\in{\mathcal{R}}$. If $(p_{u},u)\notin{\mathcal{R}}$, then the degree constraints of the vertices in $\mathcal{D}$} imply that $(u,c_{u})\notin{\mathcal{M}}$. On the other hand, if $(p_{u},u)\in{\mathcal{R}}$, then $(u,c_{u})\in{\mathcal{R}}$ and, in turn, either both of $(p_{u},u)$ and $(u,c_{u})$ are in $\mathcal{M}$} or neither. Now obtain $\mathcal{M}$$^{\prime}$ from $\mathcal{M}$} by replacing $(p_{u},u)$ and $(u,c_{u})$ with $(p_{u},c_{u})$ if $(p_{u},u)$ and $(u,c_{u})$ are in $\mathcal{M}$}, and replacing $f$ with $(p_{u^{\prime}},u^{\prime})$ and $(u^{\prime},c_{u^{\prime}})$ if $f\in{\mathcal{M}}$. Thus, as $\mathcal{M}$} is the embedding of $\mathcal{D}$} in $\mathcal{N}$} that underlies $\mathcal{R}$}, it follows from the construction of $\mathcal{N}$$^{\prime}$ from $\mathcal{N}$} that $\mathcal{M}$$^{\prime}$ is also an embedding of $\mathcal{D}$} in $\mathcal{N}$$^{\prime}$. Hence, $\mathcal{D}$} is a tree-child digraph of $\mathcal{N}$$^{\prime}$. To complete the proof, let $\mathcal{R}$$^{\prime}$ be the digraph obtained from $\mathcal{R}$} by applying the following operations. If $u$ has in-degree zero in $\mathcal{R}$}, then delete $u$ and, if $u$ has in-degree one in $\mathcal{R}$}, then suppress $u$. Moreover, if $f\notin{\mathcal{R}}$, then add $(u^{\prime},c_{u^{\prime}})$ and, if $f\in{\mathcal{R}}$, then subdivide $(p_{u^{\prime}},c_{u^{\prime}})$ with $u^{\prime}$. Again, by Lemma 5, $u$ has either in-degree zero and out-degree one, or in-degree one and out-degree one in $\mathcal{R}$} and, so, $\mathcal{R}$$^{\prime}$ is well defined. As $\mathcal{R}$} is an extension of $\mathcal{D}$} in $\mathcal{N}$} with $e\notin{\mathcal{R}}$ and $(u^{\prime},v)\notin{\mathcal{R}}^{\prime}$, it now follows that $\mathcal{R}$$^{\prime}$ is an extension of $\mathcal{D}$} in $\mathcal{N}$$^{\prime}$ with $|E_{{\mathcal{N}}}-E_{{\mathcal{R}}}|=|E_{{\mathcal{N}}^{\prime}}-E_{{\mathcal{R}}^{\prime}}|$. ∎
The next lemma and corollary show that there always exists a tree-child $\mathrm{SNPR}$ sequence connecting two tree-child networks with certain desirable properties. These properties will be leveraged later to establish one of the two inequalities of Theorem 1.
Lemma 11.
@@
    $({\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{i},{\mathcal{N}}_{i+3},{\mathcal{N}}_{i+4},\ldots,{\mathcal{N}}_{t-1},{\mathcal{N}}_{t})$ is a tree-child $\mathrm{SNPR}$ sequence connecting ${\mathcal{N}}_{0}$ and ${\mathcal{N}}_{t}$ of length t2t-2,
@@
    $({\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{i},{\mathcal{N}}_{i+2},{\mathcal{N}}_{i+3},\ldots,{\mathcal{N}}_{t-1},{\mathcal{N}}_{t})$ is a tree-child $\mathrm{SNPR}$ sequence connecting ${\mathcal{N}}_{0}$ and ${\mathcal{N}}_{t}$ of length t1t-1 such that ${\mathcal{N}}_{i+2}$ is obtained from ${\mathcal{N}}_{i}$ by an $\mathrm{SNPR}^{\pm}$,
@@
    $({\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{i},{\mathcal{N}}_{i+2},{\mathcal{N}}_{i+3},\ldots,{\mathcal{N}}_{t-1},{\mathcal{N}}_{t})$ is a tree-child $\mathrm{SNPR}$ sequence connecting ${\mathcal{N}}_{0}$ and ${\mathcal{N}}_{t}$ of length t1t-1 such that ${\mathcal{N}}_{i+2}$ is obtained from ${\mathcal{N}}_{i}$ by an $\mathrm{SNPR}^{-}$, or
@@
    there exists a tree-child $\mathrm{SNPR}$ sequence (N_0,N_1,N_2,…,N_i,N_i+1’,N_i+2,N_i+3,…,N_t-1,N_t) connecting ${\mathcal{N}}_{0}$ and ${\mathcal{N}}_{t}$ of length $t$ such that ${\mathcal{N}}_{i+1}^{\prime}$ is obtained from ${\mathcal{N}}_{i}$ by an $\mathrm{SNPR}^{-}$ and ${\mathcal{N}}_{i+2}$ is obtained from ${\mathcal{N}}_{i+1}^{\prime}$ by an $\mathrm{SNPR}^{+}$ (resp. $\mathrm{SNPR}^{\pm}$).
@@
Suppose that there exists an element $i\in\{0,1,2,\ldots,t-2\}$ such that ${\mathcal{N}}_{i+1}$ is obtained from ${\mathcal{N}}_{i}$ by an SNPR+ and ${\mathcal{N}}_{i+2}$ is obtained from ${\mathcal{N}}_{i+1}$ by an SNPR-. Let $e=(u,v)$ be the reticulation edge of ${\mathcal{N}}_{i+1}$ that is added in obtaining ${\mathcal{N}}_{i+1}$ from ${\mathcal{N}}_{i}$. Furthermore, let $e^{\prime}=(u^{\prime},v^{\prime})$ be the reticulation edge in ${\mathcal{N}}_{i+1}$ that is deleted in obtaining ${\mathcal{N}}_{i+2}$ from ${\mathcal{N}}_{i+1}$. If $e=e^{\prime}$, then ${\mathcal{N}}_{i}\cong{\mathcal{N}}_{i+2}$, and so (i) holds. We may therefore assume that $e\neq e^{\prime}$. If $v=v^{\prime}$, let $w$ be the child of $v$ in ${\mathcal{N}}_{i+1}$, and let $c_{u}$ and $p_{u}$ be the child and parent, respectively, of $u$ in ${\mathcal{N}}_{i+1}$ such that $c_{u}\neq v$. Since $u$ is a tree vertex, $c_{u}$ and $p_{u}$ are well defined. Observe that $(u^{\prime},w)$ and $(p_{u},c_{u})$ are tree edges in ${\mathcal{N}}_{i}$. It now follows that ${\mathcal{N}}_{i+2}$ can be obtained from ${\mathcal{N}}_{i}$ by the SNPR± that deletes $(u^{\prime},w)$, suppresses $u^{\prime}$, subdivides $(p_{u},c_{u})$ with a new vertex $u$, and adds the edge $(u,w)$, and so (ii) holds. So assume that $e\neq e^{\prime}$ and $v\neq v^{\prime}$. As ${\mathcal{N}}_{i+1}$ is tree-child, it follows that $e^{\prime}$ is not incident with $u$ or $v$. Thus, $e^{\prime}$ is a reticulation edge of ${\mathcal{N}}_{i}$. Let ${\mathcal{N}}_{i+1}^{\prime}$ be the phylogenetic network obtained from ${\mathcal{N}}_{i}$ by deleting $e^{\prime}$ and suppressing $u^{\prime}$ and $v^{\prime}$. By Lemma 4, ${\mathcal{N}}_{i+1}^{\prime}$ is tree-child. It is now straightforward to check that ${\mathcal{N}}_{i+2}$ can be obtained from ${\mathcal{N}}_{i+1}^{\prime}$ by an SNPR+, and so (iv) holds.
[…]
Figure 7: The three cases in the last case analysis in the proof of Lemma 11, where ${\mathcal{N}}_{i+1}$ is obtained from ${\mathcal{N}}_{i}$ by an SNPR±, ${\mathcal{N}}_{i+2}$ is obtained from ${\mathcal{N}}_{i+1}$ by an SNPR-, and $p_{u^{\prime}}\neq p_{u}$ or $c_{u^{\prime}}\neq c_{u}$. The three cases are (I) $e^{\prime}=(p_{u},c_{u})$, (II) $e^{\prime}=(u^{\prime},c_{u^{\prime}})$, and (III) $e^{\prime}\neq(p_{u},c_{u})$ and $e^{\prime}\neq(u^{\prime},c_{u^{\prime}})$. Vertex labels $y_{1}$, $y_{2}$, and $y_{3}$ are only used to clarify the figure.